OpenStudy (anonymous):

ln x + ln 15 write the expression as a single logarithm whose coefficient is 1

OpenStudy (saifoo.khan):

\[\log a + \log b = \log(a\times b)\]

OpenStudy (anonymous):

saifooooo! what s with the hair?

OpenStudy (saifoo.khan):

s??

OpenStudy (anonymous):

= ln15x?

OpenStudy (saifoo.khan):

yes

OpenStudy (anonymous):

that seems too easy! thank you